DIRC Meeting Operations Report

Period: Wednesday 2007/3/28 - Wednesday 2007/4/24

Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland, Jose Benitez Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070424

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/DIRCmeeting_20070424.html

Contents


BaBar-PEP news

PEP peak lumi: ~1.0 x 1034 (previous week: ~1.0 x 1034)

BaBar integrated lumi: 2.13 fb-1 (previous week: 3.28 fb-1)

Run 6 PEP delivery: 2.19 fb-1 (previous week: 3.43 fb-1)

Run 6 total delivery: ~30.61 fb-1 Run 6 total BaBar recorded: 29.38 fb-1 including 0 fb-1 off-peak data

PEP activities since this week:

  • Monday (4/9): VacSealing made on 4/9 near Q4.
  • Wednesday (4/18): in the morning 2 hours MD for beam characterization.
  • Thursday (4/19): 12 hours ROD.
  • Over the weekend (4/20 - 4/23): problems since skew quads installation. Frequent MD studies at low luminosity.

    BaBar activities/issues this week:

  • 4/9: 1 day ROD.
  • 4/19: 12 hours ROD.

    DIRC activities

    See DIRC logbook for details

    Backgrounds

    Scheduled activities

    Pending Work

    Spare status

    Manpower situation

    Jose, Alejandro and Bradley rotating backup & on-call commissioning duties.

    Links